Can someone please help me calculate the current wacc of Apple right now? I would also like to know step by step if possible. Thank you
Answer #1

Answer: This is the latest WACC of Apple Inc. (As on October 30, 2019)

WACC- It is the weighted average cost of capital that is the average cost of capital that a company has to pay to the security holders to finance the assets.

Formula of WACC = Re * (E/V) + Rd (1-T) * (D/V)

Where

  • Re = Cost of Equity
  • E/V = Equity percentage in the value of the firm
  • Rd = Cost of debt
  • D/V = Debt percentage in the value of the firm
  • T = Tax rate

Calculating Cost of equity from CAPM model:

Re = Rf + Beta (Rm-Rf)

Where Rf is risk free rate and (Rm- Rf) = Risk premium

Re: 1.78% + 1.10 * 6% = 8.38%

Calculating cost of debt:

Interest expenses / Last two years average debt (book value)

Cost of debt: 3240 / 115081.50 = 2.8154%

Capital structure = Equity + Debt

Equity % in Apple's capital structure = .9052

Debt % in Apple's capital structure = .0984

Average Tax rate (Latest two years) = 21.45%

Putting all the given values in the formula, we get:

WACC = .0838 * .9052 + .028154 * (1-.02145) *.0984

WACC = .07585576 + .0027109

WACC = .07856666

WACC = 7.86%